Please can you help me.
What is the value of x
((( 8x/3x + 2x = 0 )))

Answer by stanbon(48535)

8x/3x + 2x = 0
Multiply thru by 3x to get:
8x+6x^2=0
6x^2+8x=0
Factor to get:
2x(3x+4)=0
x=0 or x=-4/3
But x cannot be zero in the orinal equation.
So. x=-4/3
Cheers,
Stan H.